James is 12 years old. He will be 5/3 the age of his sister in 3 years. How old will the sister be in two years?
Answer by 100601(6)

J= age of James now
S= age of James' sister now
J=12
J+3=5/3(S+3)
12+3=5/3(S+3)
15= 5/3S+5
10= 5/3S
S= 6
6+2= 8